OpenStudy (anonymous):

Find the real solutions of the equation: 6+√(2x-9)=x

OpenStudy (anonymous):

\[6+\sqrt{2x-9}=x \implies \sqrt{2x-9}=x-6\]\[\implies 2x-9=x^2-12x+36 \implies 0=x^2-10x+45\]

OpenStudy (anonymous):

That ought to get you close enough to solve it yourself. Don't forget to check your answers to make sure they work.

OpenStudy (anonymous):

Could you give me the answer I'm stuck... Do I have to factor it?

OpenStudy (anonymous):

Actually, there are no real solutions, since b^2-4ac=100-180<0.